# Efficient Retrieval with Learned Similarities

### v2

@@ Appendix A Experiment Setups
We follow the implementation provided by the authors [42], which parameterizes $\pi_{p}(q,x)$ as a two-layer multi-layer perceptron (MLP) with SiLU [7] non-linearity. The inputs to this MLP consist of user-side features, item-side features, and the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the learned low-rank embeddings.
Hyperparameter settings.
@@
We use an identical number of sampled negatives for dot product baselines (cosine similarity, “SASRec”, “HSTU” rows in Table 3) and Mixture-of-logits (“SASRec + MoL”, “HSTU + MoL” rows in Table 3) to ensure a fair comparison, which is $128$ for ML-1M and ML-20M and $512$ for Amazon Books following prior work. For “+ MoL” rows, we additionally grid searched $|P_{x}|$ in $\{2,4,8,16\}$, $d_{P}$ in $\{32,64,128\}$), whether to enable user-id based learned embeddings, and the dropout rate to apply to user-id based embeddings in $\{0.2,0.5,0.8\}$ for the smaller MovieLens datasets. Due to resource constraints, we followed initial hyperparameters provided by the authors [42] for all other parameters.
A.1 Efficiency
Method HR@10 HR@50 MRR Time/Epoch
ML-1M SASRec .2818 .5470 .1352 37s
SASRec + MoL .3034 (+7.7%) .5655 (+3.4%) .1448 (+7.1%) 44s (+18.9%)
HSTU .3332 .5956 .1579 41s
HSTU + MoL .3460 (+3.8%) .6048 (+1.5%) .1673 (+6.0%) 46s (+12.2%)
ML-20M SASRec .2883 .5484 .1375 197s
SASRec + MoL .3124 (+8.4%) .5685 (+3.7%) .1551 (+12.8%) 449s (+127.9%)
HSTU .3557 .6146 .1800 526s
HSTU + MoL .3697 (+3.9%) .6244 (+1.6%) .1880 (+4.4%) 773s (+47.0%)
Books SASRec .0306 .0754 .0153 506s
SASRec + MoL .0402 (+31.4%) .0970 (+28.6%) .0203 (+32.7%) 964s (+90.5%)
HSTU .0469 .1066 .0233 1883s
HSTU + MoL .0613 (+30.7%) .1292 (+21.2%) .0315 (+35.2%) 2236s (+18.7%)
Table 5: Evaluation of recommendation retrieval performance, utilizing the next-item prediction task. Note that these results are identical to those reported in Table 3 but with relative gains (MoL vs dot product) and training cost changes illustrated under different sequential encoder backbones.
Training efficiency.
We measure the training efficiency of the setups in Section 4 over one epoch of data, utilizing a single RTX 6000 Ada GPU and with a constant batch size of 128 for ML-1M and ML-20M, and 64 for Amazon Books in Table 5. Consistent with prior findings [42, 3], MoL’s model quality increases do not generally come with a large increase in training time, thanks to MoL’s higher arithmetic intensity on GPUs, with an average increase of 53% across the six scenarios. To contextualize this 1.53x training cost, prior work on sequential recommenders, such as self-supervised bidirectional models, came with a 5.9x training time increase on ML-20M and 4.5x training time increase on Amazon Reviews (BERT4Rec vs SASRec rows) [23]. We expect additional GPU kernel-level optimizations, such as kernel fusions, to further improve efficiency of the “+ MoL” rows.
Inference efficiency.
Given the number of items (3K, 24K, and 674K) in these datasets, as discussed in Table 2, these public benchmark datasets are not good candidates for measuring inference latency. We note that prior work have reported batched inference latency between 5ms and 20ms on datasets with hundreds of millions to billions of items [42, 3], which is inline with optimized MIPS baselines.

### v4

@@ A.3. Parameterization of $\pi_{p}(q,x)$ matrices
We follow the implementation provided in the original MoL paper (Zhai et al., 2023), which parameterizes $\pi_{p}(q,x)$ as a two-layer multi-layer perceptron (MLP) with SiLU (Elfwing et al., 2017) non-linearity. For recommendation datasets (ML-1M, ML-20M, Books), the inputs to this MLP consist of user-side features, item-side features, and the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the low-rank embeddings. For question answering datasets (NQ320K), we only use the last part – the $P$ dot products $\langle f_{p}(q),g_{p}(x)\rangle$s between the low-rank embeddings – as inputs to this MLP.
A.4. Hyperparameter settings
A.4.1. Recommendation Systems
@@
We use an identical number of sampled negatives for dot product baselines (cosine similarity, “SASRec”, “HSTU” rows in Table 3) and Mixture-of-Logits (“SASRec + MoL”, “HSTU + MoL” rows in Table 3) to ensure a fair comparison, which is $128$ for ML-1M and ML-20M and $512$ for Amazon Books following prior work. For “+ MoL” rows, we additionally grid searched $|P_{x}|$ in $\{2,4,8,16\}$, $d_{P}$ in $\{32,64,128\}$, whether to enable user-id based learned embeddings, and the dropout rate to apply to user-id based embeddings in $\{0.2,0.5,0.8\}$ for the smaller MovieLens datasets. We followed initial hyperparameters provided by the authors (Zhai et al., 2023) for all other parameters. The models are trained using PyTorch over 1 NVIDIA RTX 6000 Ada GPU for the smaller ML-1M and ML-20M datasets and 2 NVIDIA RTX 6000 Ada GPUs for the larger Books datasets.
A.4.2. Question Answering (QA)
We train the model with AdamW optimizer (Loshchilov and Hutter, 2019), and grid searched learning rate in {2e-4, 5e-4, 8e-4} due to the introduction of the parameterized pooling component (Appendix A.2). We apply linear scheduling with warm-up over a fixed 10% of the training epochs. We train the model on 4 NVIDIA H100 80GB GPUs with a local batch size of 512. Note that due to the computational requirements of this dataset, prior work are frequently trained on 8 GPUs (Karpukhin et al., 2020; Wang et al., 2022b) or more, e.g., 32 GPUs in GENRE (De Cao et al., 2021) and 256 TPUs in DSI (Tay et al., 2022). We perform in-batch negative sampling, consistent with baselines (Karpukhin et al., 2020; Ni et al., 2022a). For MoL hyperparameters, we grid searched $P_{Q}$ and $P_{X}$ in {(2, 2), (4, 4), (8, 8), (16, 16)}, kept $d_{P}$ identical to the embedding dimension of the pretrained language model ($768$), and selected the best hyperparameters utilizing a validation set.
Appendix B Examples for exact and approximate top $K$ retrieval algorithms
We provide examples for our exact- and approximate- retrieval algorithms discussed in Section 3 to facilitate understanding.
Retrieval with Exact Algorithm (Section 3.1).
item $\langle f_{1},g_{1}\rangle$ $\langle f_{2},g_{2}\rangle$ $\pi_{1}$ $\pi_{2}$ $\phi$
a 1 1 0.5 0.5 1.0
b 0.8 0 0.5 0.5 0.4
c 0 0.8 0.5 0.5 0.4
d 0.7 0 1 0 0.7
e 0.2 0.2 0.5 0.5 0.2
Table 6. Example illustrating how exact- and approximate- top-$k$ algorithms work. We consider a fixed query $q$, and provide inner products $\langle f_{p}(x),g_{p}(q)\rangle$s, gating weights $\pi_{p}$s, and learned similarity scores $\phi$s for that query.
Table 6 shows an example with $4$ items and $2$ pairs (groups) of embeddings ($P=2$). Assume the goal is to retrieve the top $K=2$ items. In the first stage, we retrieve the top $2$ items for each embedding set, i.e., item $a,b,c$ are retrieved. For each of the retrieved items, we calculate the MoL scores based on their gating weights, i.e., $1.0,0.4,0.4$ for $a,b$, and $c$, respectively. Here, $S_{min}$ is set to $0.4$. In the second stage, we retrieve all items with $\langle f,g\rangle\geq 0.4$ for each embedding set, i.e., item $d$, and then calculate their corresponding MoL score, i.e., $0.7$. The algorithm returns $a$ and $d$ as the top $2$ items.
Retrieval with Approximate Algorithm (Theorem 3).
Consider again the example shown in Table 6. Assume we want to retrieve the top $2$ items. With top $K$ over component-level embeddings, item $a,b,c$ are retrieved, and $S_{k}=0.4$. Since $S$ is calculated as $\max\{0.7,0.2\}$, the upper bound of the gap is $0.3$. Here, the gap is exact, i.e., the actual second largest MoL score with item $d$ is $0.3$ higher than the second largest MoL score from the set of retrieved items $\{a,b,c\}$. If we bound the gap with the retrieved items only, i.e., $a,b,c$, then we will get a looser bound of $0.8-0.4=0.4$.